# The Unbearable Slowness of Being

### v1

@@ 5.1 How much information is stored in the human brain?
This is a popular question when people compare the brain to computers or consider the needs for the artificial intelligence systems of the future. It is generally thought that memory is stored in the strengths of synaptic connections, and to some extent in the properties of individual neurons. So we can obtain an upper bound on the brain’s capacity by counting the number of synapses ($\sim 10^{14}$) and multiplying by the dynamic range of synaptic strength, estimated at 5 bits [bartol_nanoconnectomic_2015], resulting in $5\times 10^{14}{\textrm{ bits}}\approx 50{\textrm{ Terabytes}}$ required to specify all synapses. One may want to add a few bits per neuron to specify its functional parameters, but given the small number of neurons ($\sim 10^{11}$), that will leave the estimate at the same order of magnitude. This calculation assumes that any combination of connections is possible and leads to a viable brain, so the upper bound of 50 TB is also a vast overestimate.
One way to tighten the bound is to consider how information gets into the brain. There are two routes: nature and nurture. The former is implemented by the biological mechanisms of early development, largely steered by a dynamic program of gene expression that unfolds approximately the same in every member of the species. The outcome of this process is encoded in the genome. The human genome contains $\sim 3\times 10^{9}$ bases. Humans differ from each other at only a tiny fraction of these sites: only a small fraction of the bases encode genes, and not all of the genes are expressed in the brain [hawrylycz_anatomically_2012]. Still, let us suppose generously that all bases in the genome independently control brain development, with each base governing a different dimension of brain connectivity. That places an upper bound on the “nature” component at $\sim 6\times 10^{9}$ bits or $\sim$ 0.8 GB. Again, this calculation pretends that every possible genome will lead to a viable brain, and thus represents a vast overestimate.
Nurture contributes to brain development through the information acquired via the senses. As we have seen, humans at their very best can take in about 10 bits/s from the environment. So even if a person soaks up information at the perceptual limit of a Speed Card champion, does this 24 hours a day without sleeping, and lives for 100 years, they will have acquired approximately $3\times 10^{10}$ bits $<4$ GB. After including the information from the genome, this still fits comfortably on the key-chain thumb drive you carry in your pocket.
@@ 5.2 The speed of life for different species
How can humans get away with just 10 bits/s? The tautological answer here is that cognition at such a low rate is sufficient for survival. More precisely, our ancestors have chosen an ecological niche where the world is slow enough to make survival possible. In fact, the 10 bits/s are needed only in worst-case situations, and most of the time our environment changes at a much more leisurely pace. This contributes to the common perception among teenagers that “reality is broken” [mcgonigal_reality_2011], leading them to seek solace in fast-paced video games (see Appendix B). Previous generations are no exceptions – they instead sought out the thrills of high-speed sports like skiing or mountain biking. It appears that pushing ourselves to the cognitive throughput limit is a rewarding experience all by itself.
@@
Remarkably there has been very little research in this area. For example, the visual system of flies has served as the proving ground for information-theoretic measurements in neuroscience [de_ruyter_van_steveninck_rate_1996, bialek_reading_1991, borst_information_1999]. Yet our survey of researchers in this area could find no attempt to measure the information throughput of the whole fly. Some relevant raw data exist. For example, theobald_dynamics_2010 measured the torque that a tethered flying fly produces in response to a horizontal rotation of the visual environment. By analyzing the published time series using the methods reviewed in borst_information_1999 we obtained an information rate of 0.62 bits/s. Note this is 100-fold smaller than the information rate about this kind of stimulus obtained from a single neuron “H1” in the fly’s visual system [bialek_reading_1991]. Granted, the torque measurements of theobald_dynamics_2010 apply to only one dimension of flight control, and a freely moving insect may operate more efficiently. Nonetheless, it seems likely that the paradoxical contrast between low behavioral throughput and high single-neuron information rates exists in insects as well.
A serious study of the overall behavioral throughput in different species could be illuminating. It is generally recognized that gathering information has fitness value [donaldson-matasci_fitness_2010, kussell_phenotypic_2005]. Vice versa information processing entails a metabolic load and thus comes at a fitness cost [laughlin_metabolic_1998, lennie_cost_2003]. But how did different lineages manage this trade-off? Does each species operate at a throughput that is just sufficient for survival? How does that depend on their ecological niche? Does the “speed of being” actually contribute to speciation into different ecological niches?
One species that operates at much higher rates is machines. Robots are allowed to play against humans in StarCraft tournaments, using the same sensory and motor interfaces, but only if artificially throttled back to a rate of actions that humans can sustain [vinyals_starcraft_2017]. It is clear that machines will excel at any task currently performed by humans, simply because their computing power doubles every two years [mckenzie_moores_2023]. So the discussion of whether autonomous cars will achieve human-level performance in traffic already seems quaint: roads, bridges, and intersections are all designed for creatures that process at 10 bits/s. When the last human driver finally retires, we can update the infrastructure for machines with cognition at kilobits/s. By that point, humans will be advised to stay out of those ecological niches, just as snails should avoid the highways.
@@ 5.3 The Musk illusion
Many people believe that their internal life is much richer than anything they can express in real time through their mouth or otherwise. One can view this illusion as a version of subjective inflation (see 4.2) generalized beyond visual perception: Because we could engage in any one of the $2^{10}$ possible actions or thoughts in the next second, it feels as though we could execute them all at the same time. In practice, however, they happen sequentially.
For the most part, this is a harmless illusion. However, when paired with the immense fortune of Elon Musk, the belief can lead to real-world consequences. Musk decided to do something about the problem and create a direct interface between his brain and a computer to communicate at his unfettered rate: “From a long-term existential standpoint, that’s, like, the purpose of Neuralink, to create a high-bandwidth interface to the brain such that we can be symbiotic with AI,” he said. “Because we have a bandwidth problem. You just can’t communicate through your fingers. It’s just too slow.” [info_clips_elon_2018]
@@ 5.4 Brain-computer interfaces
A similar argument applies to brain-computer interfaces (BCIs) intended for patients who are impaired in sensory perception or motor control. For example, in certain forms of blindness, the photoreceptors of the retina die, but the retinal ganglion cells and optic nerve fibers remain intact. One approach to restoring vision has been to implant electrode arrays in the eye and stimulate the ganglion cells there directly with signals derived from a video camera [stingl_electronic_2013, weiland_retinal_2014]. Of course, this requires transferring raw image information into the peripheral visual system, where – as reviewed above – the data rates are gigabits per second. As one might have predicted, that approach has been spectacularly unsuccessful: After decades of efforts, all the implanted patients remain legally blind [ayton_update_2020]. The major companies behind this approach have now gone out of business, and their patients are left carrying useless hardware in their eyeballs [strickland_their_2022, bioregio_stern_management_gmbh_retina_2019].
@@
At the same time, we know that humans never extract more than about 10 bits/s from the visual scene. So one could instead convey to the user only the important results of visual processing, such as the identity and location of objects and people in the scene. This can be done comfortably using natural language: A computer translates the visual scene to speech in real time and narrates it to the user according to their needs. Such a device was practical already in 2018 [liu_augmented_2018], and the intervening developments in computer vision and natural language processing are enabling even more powerful augmented-reality apps for blind people.
@@
On the motor output side, neural prostheses have been developed that aim to restore some mobility to paralyzed patients. To bridge the gap between the brain and muscles, one can implant a 100-electrode array in some region of the motor or pre-motor cortex, record the spike trains of neurons there, and decode from those signals the user’s intended movement [hochberg_reach_2012, andersen_thought_2019]. The result gets translated into commands for a robot or exoskeleton. These systems have demonstrated some interesting capabilities, for example, a recent benchmark-setting BCI can decode intended handwriting at 90 English characters per minute, or 1.5 bits/s [willett_high-performance_2021]. A speech BCI can decode intended speech up to 62 words per minute, half the rate of the typist [willett_high-performance_2023].
@@
However, the vast majority of paralyzed people are able to hear and speak, and for those patients language offers a much simpler brain-machine interface. The subject can write by simply dictating text to the machine. And she can move her exoskeleton with a few high-level voice commands. If the robot thinks along by predicting the user’s most likely requests, this communication will require only a few words (“Siri: sip beer”), leaving most of the language channel open for other uses. The important principle for both sensory and motor BCIs is that one really needs to convey only a few bits per second to and from the brain, and those can generally be carried by interfaces that don’t require drilling holes in the user’s head [chaudhary_brain-computer_2016].
@@ 6.1 Serial vs parallel processing
The vast gulf between peripheral and central information rates results in large part from the difference between parallel and serial processing. For example, the peripheral visual system performs image processing in a massively parallel fashion. The retina produces 1 million output signals, each of which is the result of a local computation on the visual image. Subsequently, the primary visual cortex takes over, using a parallel set of 10,000 modules called “hypercolumns”. Each of these covers about 0.5 mm2 of the cortical surface, contains $\sim$100,000 neurons, and elaborates a complete set of feature descriptors for a small patch of the visual field [tso_whither_2009, adams_complete_2007].
In contrast, central processing appears to be strictly serial: When faced with two tasks in competition, individuals consistently encounter a “psychological refractory period” before being able to perform the second task [broadbent_perception_1958, pashler_dual-task_1994]. Even in tasks that do not require any motor output, such as thinking, we can pursue only one strand at a time.
It helps to imagine how things could be different. For example, one of the celebrated achievements of human cognition is the “cocktail party effect”, namely our ability to extract one stream of conversation from the chaotic auditory mix in the room and to follow that speaker’s string of words [cherry_experiments_1953, bronkhorst_cocktail-party_2015]. But why do we follow just one conversation, rather than tracking all the conversations in parallel? Certainly, hundred pieces of gossip would be more useful than just one! Our peripheral auditory system operates in parallel, for example processing different frequency channels at the same time, or computing various interaural time delays. Why does this parallel processing have to end when it comes to the interesting content of audition? From this standpoint, our need to focus attention on just one conversation is a deplorable bug, not a feature.
Another useful perspective comes from the world of competitive games. Chess players decide on the next move, at least in part, based on planning. They consider various options, follow that line several moves into the future, and then evaluate the resulting positions. Why can we not evaluate these possible trajectories in parallel, all at the same time? Given the time limit set on the chess match, a player with that ability would gain a powerful advantage. Yet the psychology of chess is quite clear on the fact that even grandmasters contemplate possible moves serially, one at a time [gobet_psychology_2018].
@@ 6.2 Evolutionary history
The only living things with a nervous system are the ones that move around. Accordingly, it has been argued that the primary purpose of the brain is the control of movement [koziol_consensus_2014]. In early metazoan evolution, a simple nervous system likely served to guide its owner toward food or away from predators [budd_early_2015]. Such a primordial ocean creature might have sensory neurons on the surface by which it can detect an odor gradient. A circuit of interneurons combines this information to ultimately drive the muscles that steer and carry the animal in the desired direction. In lineages alive today, one can still recognize that the core parts of the brain are organized around an axis for olfactory navigation [jacobs_chemotaxis_2012, aboitiz_olfaction_2015].
@@
Human thinking can be seen as a form of navigation in a space of abstract concepts. It is like executing movements without activating the muscles. For example, the cerebellum – long recognized as an organ for motor control – has certain compartments dedicated to cognition [koziol_consensus_2014, mendoza_motor_2014]. These areas interact with the prefrontal cortex, instead of the motor cortex [middleton_basal_2000].
Another pointer to the primacy of spatial navigation comes from the study of memory athletes [luria_mind_1987]. Faced with the abstract task of storing long sequences of items, these performers typically imagine a navigation narrative – like a walk through a familiar neighborhood – and insert the items to be memorized into locations along that route. During the recall phase of the task, they reenter that “memory palace” and “see” the items along the route [krokos_virtual_2019]. Clearly, these elaborate spatial representations serve as an esoteric hack to fit the task of memorizing binary digits into our natural cognitive architecture.

### v2

@@ 6.1 How much information is stored in the human brain?
This is a popular question when people compare the brain to computers or consider the needs for the artificial intelligence systems of the future. It is generally thought that memory is stored in the strengths of synaptic connections, and to some extent in the properties of individual neurons. So we can obtain an upper bound on the brain’s capacity by counting the number of synapses ($\sim 10^{14}$) and multiplying by the dynamic range of synaptic strength, estimated at 5 bits bartol_nanoconnectomic_2015, resulting in $5\times 10^{14}{\textrm{ bits}}\approx 50{\textrm{ Terabytes}}$ required to specify all synapses. One may want to add a few bits per neuron to specify its functional parameters, but given the small number of neurons ($\sim 10^{11}$), that will leave the estimate at the same order of magnitude. This calculation assumes that any combination of connections is possible and leads to a viable brain, so the upper bound of 50 TB is also a vast overestimate.
One way to tighten the bound is to consider how information gets into the brain. There are two routes: nature and nurture. The former is implemented by the biological mechanisms of early development, largely steered by a dynamic program of gene expression that unfolds approximately the same in every member of the species. The outcome of this process is encoded in the genome. The human genome contains $\sim 3\times 10^{9}$ bases. Humans differ from each other at only a tiny fraction of these sites: only a small fraction of the bases encode genes, and not all of the genes are expressed in the brain hawrylycz_anatomically_2012. Still, let us suppose generously that all bases in the genome independently control brain development, with each base governing a different dimension of brain connectivity. That places an upper bound on the “nature” component at $\sim 6\times 10^{9}$ bits or $\sim$ 0.8 GB. Again, this calculation pretends that every possible genome will lead to a viable brain, and thus represents a vast overestimate.
Nurture contributes to brain development through the information acquired via the senses. As we have seen, humans at their very best can perceive and memorize about 10 bits/s from the environment. So even if a person soaks up information at the perceptual limit of a Speed Card champion, does this 24 hours a day without sleeping, and lives for 100 years, they will have acquired approximately $3\times 10^{10}$ bits $<4$ GB. After including the information from the genome, this still fits comfortably on the key-chain thumb drive you carry in your pocket.
@@ 6.2 The speed of life for different species
How can humans get away with just 10 bits/s? The tautological answer here is that cognition at such a low rate is sufficient for survival. More precisely, our ancestors have chosen an ecological niche where the world is slow enough to make survival possible. In fact, the 10 bits/s are needed only in worst-case situations, and most of the time our environment changes at a much more leisurely pace. This contributes to the common perception among teenagers that “reality is broken” mcgonigal_reality_2011, leading them to seek solace in fast-paced video games (see Appendix A). Previous generations are no exceptions – they instead sought out the thrills of high-speed sports like skiing or mountain biking. It appears that the everyday tasks feel unbearably slow for these thrill-seekers, so pushing themselves to the cognitive throughput limit is a rewarding experience all by itself.
@@
Remarkably there has been very little research in this area. For example, the visual system of flies has served as the proving ground for information-theoretic measurements in neuroscience de_ruyter_van_steveninck_rate_1996, bialek_reading_1991, borst_information_1999. Yet our survey of researchers in this area could find no attempt to measure the information throughput of the whole fly. Some relevant raw data exist. For example, Theobald et al theobald_dynamics_2010 measured the torque that a tethered flying fly produces in response to a horizontal rotation of the visual environment. By analyzing the published time series using the methods reviewed by Borst & Theunissen borst_information_1999 we obtained an information rate of 0.62 bits/s. Note this is 100-fold smaller than the information rate about this kind of stimulus obtained from a single neuron “H1” in the fly’s visual system bialek_reading_1991. Granted, the torque measurements of Theobald et al theobald_dynamics_2010 apply to only one dimension of flight control, and a freely moving insect may operate more efficiently. Nonetheless, it seems likely that the paradoxical contrast between low behavioral throughput and high single-neuron information rates exists in insects as well.
A serious study of the overall behavioral throughput in different species could be illuminating. It is generally recognized that gathering information has fitness value donaldson-matasci_fitness_2010, kussell_phenotypic_2005. Vice versa information processing entails a metabolic load and thus comes at a fitness cost laughlin_metabolic_1998, lennie_cost_2003. But how did different lineages manage this trade-off? Does each species operate at a throughput that is just sufficient for survival? How does that depend on their ecological niche? Does the “speed of being” actually contribute to speciation into different ecological niches?
One species that operates at much higher rates is machines. Robots are allowed to play against humans in StarCraft tournaments, using the same sensory and motor interfaces, but only if artificially throttled back to a rate of actions that humans can sustain vinyals_starcraft_2017. It is clear that machines will excel at any task currently performed by humans, simply because their computing power doubles every two years mckenzie_moores_2023. So the discussion of whether autonomous cars will achieve human-level performance in traffic already seems quaint: roads, bridges, and intersections are all designed for creatures that process at 10 bits/s. When the last human driver finally retires, we can update the infrastructure for machines with cognition at kilobits/s. By that point, humans will be advised to stay out of those ecological niches, just as snails should avoid the highways.
@@ 6.3 The Musk illusion
Many people believe that their internal life is much richer than anything they can express in real time through their mouth or otherwise. One can view this illusion as a version of subjective inflation (see 5.2) generalized beyond visual perception: Because we could engage in any one of the $2^{10}$ possible actions or thoughts in the next second, it feels as though we could execute them all at the same time. In practice, however, they happen sequentially.
For the most part, this is a harmless illusion. However, when paired with the immense fortune of Elon Musk, the belief can lead to real-world consequences. Musk decided to do something about the problem and create a direct interface between his brain and a computer to communicate at his unfettered rate: “From a long-term existential standpoint, that’s, like, the purpose of Neuralink, to create a high-bandwidth interface to the brain such that we can be symbiotic with AI,” he said. “Because we have a bandwidth problem. You just can’t communicate through your fingers. It’s just too slow.” info_clips_elon_2018
@@ 6.4 Brain-computer interfaces
A similar argument applies to brain-computer interfaces (BCIs) intended for patients who are impaired in sensory perception or motor control. For example, in certain forms of blindness, the photoreceptors of the retina die, but the retinal ganglion cells and optic nerve fibers remain intact. One approach to restoring vision has been to implant electrode arrays in the eye and stimulate the ganglion cells there directly with signals derived from a video camera stingl_electronic_2013, weiland_retinal_2014. Of course, this requires transferring raw image information into the peripheral visual system, where – as reviewed above – the data rates are gigabits per second. While driven by good intentions, this approach has been altogether unsuccessful: After decades of efforts, all the implanted patients remain legally blind ayton_update_2020. The major companies behind this approach have now gone out of business, and their patients are left carrying abandoned hardware in their eyeballs strickland_their_2022, bioregiosternmanagementgmbh_retina_2019.
@@
At the same time, we know that humans never extract more than about 10 bits/s from the visual scene. So one could instead convey to the user only the important results of visual processing, such as the identity and location of objects and people in the scene. This can be done comfortably using natural language: A computer translates the visual scene to speech in real time and narrates it to the user according to their needs. Such a device was practical already in 2018 liu_augmented_2018, and the intervening developments in computer vision and natural language processing are enabling even more powerful augmented-reality apps for blind people.
@@
On the motor output side, neural prostheses have been developed that aim to restore some mobility to paralyzed patients. To bridge the gap between the brain and muscles, one can implant a 100-electrode array in some region of the motor or pre-motor cortex, record the spike trains of neurons there, and decode from those signals the user’s intended movement hochberg_reach_2012, andersen_thought_2019. The result gets translated into commands for a robot or exoskeleton. These systems have demonstrated some interesting capabilities, for example, a recent benchmark-setting BCI can decode intended handwriting at 90 English characters per minute, or 1.5 bits/s willett_high-performance_2021. A speech BCI can decode intended speech up to 62 words per minute, half the rate of the typist willett_high-performance_2023. Such devices could greatly benefit those with conditions like amyotrophic lateral sclerosis (ALS) and Parkinson’s disease, who struggle with speech production.
@@
However, the vast majority of paralyzed people are able to hear and speak, and for those patients language offers a much simpler brain-machine interface. The subject can write by simply dictating text to the machine. And she can move her exoskeleton with a few high-level voice commands. If the robot thinks along by predicting the user’s most likely requests, this communication will require only a few words (“Siri: sip beer”), leaving most of the language channel open for other uses. The important principle for both sensory and motor BCIs is that one really needs to convey only a few bits per second to and from the brain, and those can generally be carried by interfaces that don’t require drilling holes in the user’s head chaudhary_brain-computer_2016.
@@ 7 Constraints on the speed of cognition
7.1 Inefficient neural hardware
A frequently cited argument is that neurons are simply inefficient devices for information processing, so nature needs to employ huge numbers of them to complete even simple operations. In particular, individual neurons – unlike transistors – are subject to random biochemical noise, so perhaps one needs to average over many neurons to get a reliable signal. Also, perhaps the brain includes a large amount of redundancy, with populations of essentially identical neurons, to guard against the loss of neurons from aging or injury. Neither of these explanations seems plausible.
Regarding “noisy neurons”, the analysis reviewed above (section 4) directly addresses that point and shows that a single neuron can transmit several bits per spike about its dendritic inputs. Certain irreducible noise sources– such as thermal fluctuations of ion channels or vesicle fusion events – may well limit the rate of information transfer, but the precision of single neurons remains remarkable. A single optic nerve fiber can reliably signal the arrival of its preferred feature in the visual stimulus with a precision of 1 millisecond berry_structure_1997. Even deep in the primate cortex, a single neuron can distinguish between two visual stimuli as precisely as the entire monkey britten_analysis_1992. What was once considered irreducible noise at the population level may, in fact, be a reliable signal about variables or inputs that the experimenter cannot control ringach_spontaneous_2009, stringer_spontaneous_2019, such as an itch on the monkey’s head.
On the subject of redundancy, again there is precious little evidence for duplication of essentially identical neurons. For example, in the fruit fly, many neuron types exist in just two copies, one on each side of the brain sterne_classification_2021, aso_mushroom_2014. One might argue that fruit flies are cheap and short-lived, so nature simply duplicates the entire organism to ensure robustness. However, the principle applies even to large animals like ourselves. In the primate retina, every point in the visual field is covered by just one neuron of each cell type, with very little overlap sanes_types_2015, wassle_cone_2009, reese_design_2015. In the visual cortex, a small stroke can abolish vision in a section of the visual field. Apparently, there is no redundancy built in for one of the most common forms of brain injury.
@@ 7.2 Serial vs parallel processing
The vast gulf between peripheral and central information rates results in large part from the difference between parallel and serial processing. For example, the peripheral visual system performs image processing in a massively parallel fashion. The retina produces 1 million output signals, each of which is the result of a local computation on the visual image. Subsequently, the primary visual cortex takes over, using a parallel set of 10,000 modules called “hypercolumns”. Each of these covers about 0.5 mm2 of the cortical surface, contains $\sim$100,000 neurons, and elaborates a complete set of feature descriptors for a small patch of the visual field tso_whither_2009, adams_complete_2007.
In contrast, central processing appears to be strictly serial: When faced with two tasks in competition, individuals consistently encounter a “psychological refractory period” before being able to perform the second task broadbent_perception_1958, pashler_dual-task_1994. Even in tasks that do not require any motor output, such as thinking, we can pursue only one strand at a time.
It helps to imagine how things could be different. For example, one of the celebrated achievements of human cognition is the “cocktail party effect”, namely our ability to extract one stream of conversation from the chaotic auditory mix in the room and to follow that speaker’s string of words cherry_experiments_1953, bronkhorst_cocktail-party_2015. But why do we follow just one conversation, rather than tracking all the conversations in parallel? Certainly, hundred pieces of gossip would be more useful than just one! Our peripheral auditory system operates in parallel, for example processing different frequency channels at the same time, or computing various interaural time delays. Yet the selection of a single speaker occurs at a fairly early level, before any word segmentation of the auditory stream mesgarani_selective_2012, brodbeck_rapid_2018. Why does parallel processing have to end when it comes to the interesting content of audition? From this standpoint, our need to focus attention on just one conversation is a deplorable bug, not a feature.
Another useful perspective comes from the world of competitive games. Chess players decide on the next move, at least in part, based on planning. They consider various options, follow that line several moves into the future, and then evaluate the resulting positions. Why can we not evaluate these possible trajectories in parallel, all at the same time? Given the time limit set on the chess match, a player with that ability would gain a powerful advantage. Yet the psychology of chess is quite clear on the fact that even grandmasters contemplate possible moves serially, one at a time gobet_psychology_2018.
@@ 7.3 Evolutionary history
The only living things with a nervous system are the ones that move around. Accordingly, it has been argued for some time that the primary purpose of the brain is the control of movement anderson_brain_2016. In early metazoan evolution, a simple nervous system likely served to guide its owner toward food or away from predators budd_early_2015. Such a primordial ocean creature might have sensory neurons on the surface by which it can detect an odor gradient. A circuit of interneurons combines this information to ultimately drive the muscles that steer and carry the animal in the desired direction. In lineages alive today, one can still recognize that the core parts of the brain are organized around an axis for olfactory navigation jacobs_chemotaxis_2012, aboitiz_olfaction_2015.
@@
Human thinking can be seen as a form of navigation in a space of abstract concepts. It is like executing movements without activating the muscles. For example, the cerebellum – long recognized as an organ for motor control – has certain compartments dedicated to cognition koziol_consensus_2014, mendoza_motor_2014. These areas interact with the prefrontal cortex, instead of the motor cortex middleton_basal_2000.
Another pointer to the primacy of spatial navigation comes from the study of memory athletes luria_mind_1987. Faced with the abstract task of storing long sequences of items, these performers typically imagine a navigation narrative – like a walk through a familiar neighborhood – and insert the items to be memorized into locations along that route. During the recall phase of the task, they reenter that “memory palace” and “see” the items along the route krokos_virtual_2019. Clearly, these elaborate spatial representations serve as an esoteric hack to fit the task of memorizing binary digits into our natural cognitive architecture.